# APEX2: Adaptive and Extreme Summarization for Personalized Knowledge Graphs

### v1

@@ Abstract.
Knowledge graphs (KGs), which store an extensive number of relational facts, serve various applications. Recently, personalized knowledge graphs (PKGs) have emerged as a solution to optimize storage costs by customizing their content to align with users’ specific interests within particular domains. In the real world, on one hand, user queries and their underlying interests are inherently evolving, requiring PKGs to adapt continuously; on the other hand, the summarization is constantly expected to be as small as possible in terms of storage cost. However, the existing PKG summarization methods implicitly assume that the user’s interests are constant and do not shift. Furthermore, when the size constraint of PKG is extremely small, the existing methods cannot distinguish which facts are more of immediate interest and guarantee the utility of the summarized PKG. To address these limitations, we propose APEX2, a highly scalable PKG summarization framework designed with robust theoretical guarantees to excel in adaptive summarization tasks with extremely small size constraints. To be specific, after constructing an initial PKG, APEX2 continuously tracks the interest shift and adjusts the previous summary. We evaluate APEX2 under an evolving query setting on benchmark KGs containing up to 12 million triples, summarizing with compression ratios $\leq 0.1\%$. The experiments show that APEX outperforms state-of-the-art baselines in terms of both query-answering accuracy and efficiency.
@@ 2. Problem Definition
Query Log. Since most real queries to KGs consist of only one or two triples (Bonifati et al., 2020), assuming we have the full access to the whole KG, we study simple queries with known answers. A query log $\mathcal{Q}$ consists of a number of queries. Each query $q$ consists of a query entity $e$, query relation $r$, and a set of answer entities $\mathcal{A}$. A triple $x_{ijk}=(e_{i},r_{k},e_{j})$ in the knowledge graph $\mathcal{G}$ is said to be an answer triple to a query q, if $e_{i}=e\wedge r_{k}=r\wedge e_{j}\in\mathcal{A}$. A query may have multiple answer triples. For example, for the query “What movie did Christopher Nolan direct”, the query entity is “Christopher Nolan” (”Nolan” for short), and the query relation is “directed_movie” (”d_m” for short). Suppose in the KG all triples with the head entity “Nolan” and the relation “d_m” are (“Nolan”, “d_m”, “Interstellar”), (“Nolan”, “d_m”, “Tenet”) and (“Nolan”, “d_m”, “Oppenheimer”). Then these three triples are answer triples to the query, and the set {“Interstellar”, “Tenet”, “Oppenheimer”} is the set of answer entities. In this paper, “for triples in $q$” iterates the answer triples of $q$; “$e\in q$” iterates all entities accessed by $q$ as the query entity or answer entities; ”$\mathcal{Q}^{(t+1)}\setminus\mathcal{Q}^{(t)}$” stands for all queries in $\mathcal{Q}^{(t+1)}$ but not in $\mathcal{Q}^{(t)}$.
@@
[…]
@@
a sequence of personalized knowledge graph $\{\mathcal{P}^{(t)}\}$ of $\mathcal{G}$ for $t\in\{0,1,\ldots,T-1\}$; each $\mathcal{P}^{(t)}$, whose number of triples —$\mathcal{P}^{(t)}$— is not greater than $K$, is able to answer as many queries in $\mathcal{Q}^{(t+1)}\setminus\mathcal{Q}^{(t)}$ as possible and as correctly as possible:
@@ 3.1. Dynamic Model of User Interests
In order to formulate more conveniently, we start by freezing at a specific timestamp $T$ ($T$ as a constant), and define $\bm{q}_{total}$ to store the number of times each entity is accessed as query entity or answer entity in the query log. If accessed as the answer to a query, the marginal value will be weighted by 1# of $\frac{1}{\#\textit{ of answers}}$. We provide a simple example here. Suppose we have 5 entities (indexed 0, 1, 2, 3, 4). The first query is (entity 0, some relation, {entity 1, entity 3}), the second query is (entity 2, some relation, {entity 0, entity 3}), then $\mathbf{q}_{\rm total}$ will be a column vector $(1+0.5,0.5+0,0+1,0.5+0.5,0+0)=(1.5,0.5,1,1,0)$. Assuming the user’s temporal query log is $\mathcal{Q}^{(t)}$, we use $\mathcal{Q}^{(t)}\setminus\mathcal{Q}^{(t-1)}$ to denote new queries arriving at time $t$. Additionally, $\mathcal{Q}^{-1}=\emptyset$. Then,
@@
where for each query $i\in\mathcal{Q}^{(t)}\setminus\mathcal{Q}^{(t-1)}$ with answer set $\mathcal{A}_{i}$, $\bm{q}_{i}$ is a vector with dimension (||×$(|\mathcal{E}|\times 1)$ whose entries are
(4) $\mathbf{q}_{i}[e]=\begin{cases}1&e\ \text{is the query entity of}\ \text{query}~{}i\\ \frac{1}{|\mathcal{A}_{i}|}&e\ \text{is an answer entity to}\ \text{query}~{}i\\ 0&\text{otherwise}\ \text{(}e\ \text{is unrelated to}\ \text{query}~{}i\text{)}\end{cases}$
@@
where $r_{k}$ is the $k$-th relation in KG $\mathcal{G}$, and $\mathbf{r}$ is a vector with dimension (||×$(|\mathcal{R}|\times 1)$ as follows
@@
where $\mathbf{\tilde{q}}_{i}$ is the one-hot vector with dimension (||×$(|\mathcal{R}|\times 1)$ corresponding to the searched relation in query $i\in\mathcal{Q}^{(t)}\setminus\mathcal{Q}^{(t-1)}$ at time $t$,
@@ 3.2. Incremental Updating
An advantage of our model is that the user’s preference on entities and relations at time $T$ can be incrementally updated from the previous timestamp $T-1$. Denoting $\bm{q}^{(T)}=\sum_{i\in\mathcal{Q}^{(T)}\setminus\mathcal{Q}^{(T-1)}}\mathbf{q}_{i}$, we derive the incremental updating equations for $\mathbf{q}_{\rm total}^{(T)}$, $\mathbf{e}^{(T)}$, and $\mathbf{r}^{(T)}$ as follows.
@@
The updating of $\mathbf{H}^{(T)}$ from $\mathbf{H}^{(T-1)}$ is per-entry conditional. Assume at timestamp $T$, compared to timestamp $T-1$, if entries $i,k$ in $\mathbf{e}^{(T)}$ and entry $j$ in $\mathbf{r}^{(T)}$ are not updated (excluding decay), then $\mathbf{H}^{(T)}[i][j][k]=\gamma^{3}\mathbf{H}^{(T-1)}[i][j][k]$. Only entries in $\mathbf{H}^{(T)}[i][j][k]$ with any of those being updated need to be recalculated. The number of updated entries is small as we assign $d$ much less than the diameter of $\mathcal{G}$. Therefore, for each timestamp, we multiply $\mathbf{H}$ by $\gamma^{3}$ and then do a small-scale update.
@@ 4.1. APEX2 Framework
The incremental time complexity for APEX2’s adapting phase is $O(c\cdot|\mathcal{Q}|^{2}\cdot\log_{2}(c\cdot|\mathcal{Q}|))$, where $Q$ is the query log decomposed into sub-queries (each query in $Q$ has only one query entity, one query relation and one answer). $c=\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}$, where $nnz$ is the operator outputting the number of non-zero elements in a matrix, $\mathbf{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix of $\mathcal{G}$, $\mathcal{E}$ is the set of entities of $\mathcal{G}$. (Proof in Appendix E.2)
If set a threshold $\epsilon_{ths}$ to eliminate small-enough entries to 00, then after $log_{\gamma}\epsilon_{ths}$ timestamps, entries introduced by previous queries will be decayed to 0. In this case, the effective number of queries $|\mathcal{Q}|$ above can be bounded by a constant $log_{\gamma}\epsilon_{ths}$. By this operation, the incremental time complexity is further optimized to $O(c\cdot\log_{2}(c))$, where $c=\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}$ is the average number of neighbors within $d$-hops. In other words, the time complexity of APEX2 updating is only related to the connectivity of KG.
@@ 4.2. APEX2 Variant: APEX2-N
APEX2-N only incrementally tracks and sorts the heat of entities but not for relations. In other words, APEX2-N gives weight $1$ to entities and 00 to relations. APEX2-N is designed mainly for adaptive solutions of PKG summarization, and we leave the trade-off between weights on entities and relations to future work. Since APEX2-N is a variation of APEX2, we summarize the detailed operations of APEX2-N in Algorithm 5 in the Appendix. We also give proof of the effectiveness and efficiency of APEX2-N as follows.
@@
The incremental time complexity for APEX2-N’s adapting phase is $O(c\log_{2}(c|\mathcal{Q}|))$, where $\mathcal{Q}$ is the query log decomposed into sub-queries (each query in $\mathcal{Q}$ has only one query entity, one query relation and one answer). And $c=\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}$, where $\mathbf{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix of $\mathcal{G}$, $\mathcal{E}$ is the set of entity of $\mathcal{G}$, and $nnz$ is the operator outputting the number of non-zero elements in a matrix. (Proof in Appendix E.3)
@@ 5.1.1. Datasets
KG $||\mathcal{E}|$$||\mathcal{R}|$$|\mathcal{T}|$
@@ 5.2. Experimental Settings
Generate user queries. To calculate the average and standard deviation, we simulate 10 users to query the KGs. Following the norm set by Freebase discussed in section 3.1, we model the abstract concept of topic by ”queries with the same query entity” [^4] In fact, if a topic is defined in terms of entities, then the ability to adapt entity shift is a sufficient condition for adapting topic shift.. To simulate a real-world querying scenario with interest shift, for each KG, we generate 200 queries on 20 topics for each user. Each group of 10 consecutive queries are in the same topic. We associate each query $t$ with timestamp $t-1$. For MetaQA, we first categorize the provided queries into different topics by query entity, then randomly sample 20 distinct query entities. After that, we randomly choose 10 queries on each of the 20 topics. For DBPedia, YAGO and Freebase, we synthetically generate queries by randomly choosing 20 query entities in the KG. Then for each query, we randomly choose 10 relations (with possible multiplicity) that the query entity has. Finally, we include all entities $e$ satisfying (query entity, chosen relation, $e$) $\in\mathcal{T}$ into the answer set. We study 1-hop simple queries with known answers because in real life there is only a small portion of complex queries (Bonifati et al., 2020), which can be decomposed into simple queries.
@@ 5.3. Comparisons
Table 3. Efficiency Comparison ()($\downarrow$)( ↓ ) (unit: seconds)
@@ 5.4.2. Component Ablation
Table 4. APEX2 Ablations’ Efficiency ()($\downarrow$)( ↓ ) in seconds
@@ 5.7.3. Results
The PKG results are shown in Figure 12, 13, 14, 15. The summarized PKG at timestamp $t-1$ have been fed query $t$. (i) The sub-figure (a) (i.e., the PKGs after the first three groups of queries) of GLIMPSE and PageRank still contains all information about the out-of-interest topic ”Blue Blood” and ”Stevan Riley”, which were accessed in the first 10 queries. (ii) For both APEX2 and APEX2-N, after the first query on the new topic, the three answer triples (Chad Michael Murray, movie_to_actor, A Cinderella Story), (Chad Michael Murray, movie_to_actor, House of Wax), and (Chad Michael Murray, movie_to_actor, Left Behind) get into the PKG. However, after three times querying on the new topic, the PKG summarized by GLIMPSE only contains (Chad Michael Murray, movie_to_actor, House of Wax), and the PKG summarized by PageRank does not contain any of them. (iii) From APEX2 and APEX2-N (b) to (d), as topic on ”Chad Michael Murray” is queried again and again, the connected group centered at ”Chad Michael Murray” grows larger. At timestamp 32 (Figure d), some of its 2-hop neighbors are included, such as (House of Wax, actor_to_movie, Nicolas Cage) in APEX2 and (A Cinderella Story, director_to_movie, Mark Rosman) in APEX2-N. The reason why the actor relationship is included first is that, this relation is recently queried many times and has a high relational interest. (iv) There are some triples that are not very related to the user queries but are summarized in the GLIMPSE PKG, for example (Onibi, movie_to_language, Japanese), (john lithgow, tag_to_movie, 2010) and (Hercules, movie_to_genre, Animation). The PageRank PKG almost remains the same given three queries on the new topic. Compared to these, APEX2 and APEX2-N produce PKGs that have intuitively higher quality in terms of user’s interest.

### v2

@@ Abstract.
Knowledge graphs (KGs), which store an extensive number of relational facts, serve various applications. Recently, personalized knowledge graphs (PKGs) have emerged as a solution to optimize storage costs by customizing their content to align with users’ specific interests within particular domains. In the real world, on the one hand, user queries and their underlying interests are inherently evolving, requiring PKGs to adapt continuously; on the other hand, the summarization is constantly expected to be as small as possible in terms of storage cost. However, the existing PKG summarization methods implicitly assume that the user’s interests are constant and do not shift. Furthermore, when the size constraint of PKG is extremely small, the existing methods cannot distinguish which facts are more of immediate interest and guarantee the utility of the summarized PKG. To address these limitations, we propose APEX2, a highly scalable PKG summarization framework designed with robust theoretical guarantees to excel in adaptive summarization tasks with extremely small size constraints. To be specific, after constructing an initial PKG, APEX2 continuously tracks the interest shift and adjusts the previous summary. We evaluate APEX2 under an evolving query setting on benchmark KGs containing up to 12 million triples, summarizing with compression ratios $\leq 0.1\%$. The experiments show that APEX outperforms state-of-the-art baselines in terms of both query-answering accuracy and efficiency.
@@ 2. Problem Definition
Query Log. Since most real queries to KGs consist of only one or two triples (Bonifati et al., 2020), assuming we have the full access to the whole KG, we study simple queries with known answers. A query log $\mathcal{Q}$ consists of a number of queries. Each query $q$ consists of a query entity $e$, query relation $r$, and a set of answer entities $\mathcal{A}$. A triple $x_{ijk}=(e_{i},r_{k},e_{j})$ in the knowledge graph $\mathcal{G}$ is said to be an answer triple to a query q, if $e_{i}=e\wedge r_{k}=r\wedge e_{j}\in\mathcal{A}$. A query may have multiple answer triples. For example, for the query “What movie did Christopher Nolan direct”, the query entity is “Christopher Nolan” (”Nolan” for short), and the query relation is “directed_movie” (”d_m” for short). Suppose in the KG all triples with the head entity “Nolan” and the relation “d_m” are (“Nolan”, “d_m”, “Interstellar”), (“Nolan”, “d_m”, “Tenet”) and (“Nolan”, “d_m”, “Oppenheimer”). Then these three triples are answer triples to the query, and the set {“Interstellar”, “Tenet”, “Oppenheimer”} is the set of answer entities. In this paper, “for triples in qq” iterates the answer triples of $q$; $e\in q$” iterates all entities accessed by $q$ as the query entity or answer entities; $\mathcal{Q}^{(t+1)}\setminus\mathcal{Q}^{(t)}$” stands for all queries in $\mathcal{Q}^{(t+1)}$ but not in $\mathcal{Q}^{(t)}$.
@@
Adaptive Personalized Knowledge Graph Summarization. $\{\}$” means “a sequence of” here. $F1$ score is a conventional measure of seaching accuracy, defined in Appendix C.1.
@@
a sequence of personalized knowledge graph $\{\mathcal{P}^{(t)}\}$ of $\mathcal{G}$ for $t\in\{0,1,\ldots,T-1\}$; each $\mathcal{P}^{(t)}$, whose number of triples $\mathcal{P}^{(t)}$— is not greater than $K$, is able to answer as many queries in $\mathcal{Q}^{(t+1)}\setminus\mathcal{Q}^{(t)}$ as possible and as correctly as possible:
@@ 3.1. Dynamic Model of User Interests
In order to formulate more conveniently, we start by freezing at a specific timestamp $T$ ($T$ as a constant), and define $\bm{q}_{total}$ to store the number of times each entity is accessed as query entity or answer entity in the query log. If accessed as the answer to a query, the marginal value will be weighted by $\frac{1}{\#\textit{ of answers}}$. We provide a simple example here. Suppose we have 5 entities (indexed 0, 1, 2, 3, 4). The first query is (entity 0, some relation, {entity 1, entity 3}), the second query is (entity 2, some relation, {entity 0, entity 3}), then $\mathbf{q}_{\rm total}$ will be a column vector (1+0.5,0.5+0,0+1,0.5+0.5,0+0)=(1.5,0.5,1,1,0)(1+0.5,0.5+0,0+1,0.5+0.5,0+0)=(1.5,0.5,1,1,0). Assuming the user’s temporal query log is $\mathcal{Q}^{(t)}$, we use $\mathcal{Q}^{(t)}\setminus\mathcal{Q}^{(t-1)}$ to denote new queries arriving at time $t$. Additionally, $\mathcal{Q}^{-1}=\emptyset$. Then,
@@
where for each query $i\in\mathcal{Q}^{(t)}\setminus\mathcal{Q}^{(t-1)}$ with answer set $\mathcal{A}_{i}$, $\bm{q}_{i}$ is a vector with dimension $(|\mathcal{E}|\times 1)$ whose entries are
[…]
@@
where $r_{k}$ is the $k$-th relation in KG $\mathcal{G}$, and $\mathbf{r}$ is a vector with dimension $(|\mathcal{R}|\times 1)$ as follows
@@
where $\mathbf{\tilde{q}}_{i}$ is the one-hot vector with dimension $(|\mathcal{R}|\times 1)$ corresponding to the searched relation in query $i\in\mathcal{Q}^{(t)}\setminus\mathcal{Q}^{(t-1)}$ at time $t$,
@@ 3.2. Incremental Updating
An advantage of our model is that the user’s preference on entities and relations at time $T$ can be incrementally updated from the previous timestamp T1T-1. Denoting $\bm{q}^{(T)}=\sum_{i\in\mathcal{Q}^{(T)}\setminus\mathcal{Q}^{(T-1)}}\mathbf{q}_{i}$, we derive the incremental updating equations for $\mathbf{q}_{\rm total}^{(T)}$, $\mathbf{e}^{(T)}$, and $\mathbf{r}^{(T)}$ as follows.
@@
The updating of $\mathbf{H}^{(T)}$ from $\mathbf{H}^{(T-1)}$ is per-entry conditional. Assume at timestamp $T$, compared to timestamp T1T-1, if entries $i,k$ in $\mathbf{e}^{(T)}$ and entry $j$ in $\mathbf{r}^{(T)}$ are not updated (excluding decay), then $\mathbf{H}^{(T)}[i][j][k]=\gamma^{3}\mathbf{H}^{(T-1)}[i][j][k]$. Only entries in $\mathbf{H}^{(T)}[i][j][k]$ with any of those being updated need to be recalculated. The number of updated entries is small as we assign $d$ much less than the diameter of $\mathcal{G}$. Therefore, for each timestamp, we multiply $\mathbf{H}$ by $\gamma^{3}$ and then do a small-scale update.
@@ 4.1. APEX2 Framework
The incremental time complexity for APEX2’s adapting phase is $O(c\cdot|\mathcal{Q}|^{2}\cdot\log_{2}(c\cdot|\mathcal{Q}|))$, where $Q$ is the query log decomposed into sub-queries (each query in $Q$ has only one query entity, one query relation and one answer). $c=\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}$, where nnznnz is the operator outputting the number of non-zero elements in a matrix, $\mathbf{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix of $\mathcal{G}$, $\mathcal{E}$ is the set of entities of $\mathcal{G}$. (Proof in Appendix E.2)
If set a threshold $\epsilon_{ths}$ to eliminate small-enough entries to 0, then after $log_{\gamma}\epsilon_{ths}$ timestamps, entries introduced by previous queries will be decayed to 0. In this case, the effective number of queries $|\mathcal{Q}|$ above can be bounded by a constant $log_{\gamma}\epsilon_{ths}$. By this operation, the incremental time complexity is further optimized to $O(c\cdot\log_{2}(c))$, where $c=\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}$ is the average number of neighbors within $d$-hops. In other words, the time complexity of APEX2 updating is only related to the connectivity of KG.
@@ 4.2. APEX2 Variant: APEX2-N
APEX2-N only incrementally tracks and sorts the heat of entities but not for relations. In other words, APEX2-N gives weight $1$ to entities and 0 to relations. APEX2-N is designed mainly for adaptive solutions of PKG summarization, and we leave the trade-off between weights on entities and relations to future work. Since APEX2-N is a variation of APEX2, we summarize the detailed operations of APEX2-N in Algorithm 5 in the Appendix. We also give proof of the effectiveness and efficiency of APEX2-N as follows.
@@
The incremental time complexity for APEX2-N’s adapting phase is $O(c\log_{2}(c|\mathcal{Q}|))$, where $\mathcal{Q}$ is the query log decomposed into sub-queries (each query in $\mathcal{Q}$ has only one query entity, one query relation and one answer). And $c=\frac{nnz(\sum_{l=0}^{d}(\alpha\mathbf{A})^{l})}{|\mathcal{E}|}$, where $\mathbf{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix of $\mathcal{G}$, $\mathcal{E}$ is the set of entity of $\mathcal{G}$, and nnznnz is the operator outputting the number of non-zero elements in a matrix. (Proof in Appendix E.3)
@@ 5.1.1. Datasets
KG $\mathcal{E}| ||$$\mathcal{R}|$ $|\mathcal{T}|$
@@ 5.2. Experimental Settings
Generate user queries. To calculate the average and standard deviation, we simulate 10 users to query the KGs. Following the norm set by Freebase discussed in section 3.1, we model the abstract concept of topic by ”queries with the same query entity” [^4] In fact, if a topic is defined in terms of entities, then the ability to adapt entity shift is a sufficient condition for adapting topic shift.. To simulate a real-world querying scenario with interest shift, for each KG, we generate 200 queries on 20 topics for each user. Each group of 10 consecutive queries are in the same topic. We associate each query $t$ with timestamp t1t-1. For MetaQA, we first categorize the provided queries into different topics by query entity, then randomly sample 20 distinct query entities. After that, we randomly choose 10 queries on each of the 20 topics. For DBPedia, YAGO and Freebase, we synthetically generate queries by randomly choosing 20 query entities in the KG. Then for each query, we randomly choose 10 relations (with possible multiplicity) that the query entity has. Finally, we include all entities $e$ satisfying (query entity, chosen relation, $e$) $\in\mathcal{T}$ into the answer set. We study 1-hop simple queries with known answers because in real life there is only a small portion of complex queries (Bonifati et al., 2020), which can be decomposed into simple queries.
@@ 5.3. Comparisons
Table 3. Efficiency Comparison $(\downarrow)$ (unit: seconds)
@@ 5.4.2. Component Ablation
Table 4. APEX2 Ablations’ Efficiency $(\downarrow)$ in seconds
@@ 5.7.3. Results
The PKG results are shown in Figure 12, 13, 14, 15. The summarized PKG at timestamp t1t-1 have been fed query $t$. (i) The sub-figure (a) (i.e., the PKGs after the first three groups of queries) of GLIMPSE and PageRank still contains all information about the out-of-interest topic ”Blue Blood” and ”Stevan Riley”, which were accessed in the first 10 queries. (ii) For both APEX2 and APEX2-N, after the first query on the new topic, the three answer triples (Chad Michael Murray, movie_to_actor, A Cinderella Story), (Chad Michael Murray, movie_to_actor, House of Wax), and (Chad Michael Murray, movie_to_actor, Left Behind) get into the PKG. However, after three times querying on the new topic, the PKG summarized by GLIMPSE only contains (Chad Michael Murray, movie_to_actor, House of Wax), and the PKG summarized by PageRank does not contain any of them. (iii) From APEX2 and APEX2-N (b) to (d), as topic on ”Chad Michael Murray” is queried again and again, the connected group centered at ”Chad Michael Murray” grows larger. At timestamp 32 (Figure d), some of its 2-hop neighbors are included, such as (House of Wax, actor_to_movie, Nicolas Cage) in APEX2 and (A Cinderella Story, director_to_movie, Mark Rosman) in APEX2-N. The reason why the actor relationship is included first is that, this relation is recently queried many times and has a high relational interest. (iv) There are some triples that are not very related to the user queries but are summarized in the GLIMPSE PKG, for example (Onibi, movie_to_language, Japanese), (john lithgow, tag_to_movie, 2010) and (Hercules, movie_to_genre, Animation). The PageRank PKG almost remains the same given three queries on the new topic. Compared to these, APEX2 and APEX2-N produce PKGs that have intuitively higher quality in terms of user’s interest.